OpenStudy (anonymous):

The sum of three consecutive odd integers is –15. What are the numbers? A.{–3, –1, 1} B.{–3, –5, –7} C.{–7, –9, –11} D.{–4, –5, –6}

OpenStudy (alexandervonhumboldt2):

a+b+c=-15 a=a b=a+c c=b+1

OpenStudy (alexandervonhumboldt2):

woops b=a+1

OpenStudy (alexandervonhumboldt2):

a=a b=a+1 c=a+2

OpenStudy (alexandervonhumboldt2):

if you'l find `a`, you'll find all other numbers

OpenStudy (alexandervonhumboldt2):

woppers

OpenStudy (alexandervonhumboldt2):

a+a+1+a+2=-15

OpenStudy (alexandervonhumboldt2):

3a=-15-3 3a=-18 a=?

OpenStudy (anonymous):

It's D

OpenStudy (alexandervonhumboldt2):

yes because -4-5-6=-15

OpenStudy (anonymous):

:) Thank youu